Company Earnings
per Share
Market Value
per Share
1 $ 10.00 $

160.00

2 8.00

79.20

3 6.00

78.00

4 45.00

297.00

  
Compute the price-earnings ratio for each of these four separate companies.
Answer #1

Price-Earnings Ratio = Market Price per share/Earnings per share

Company 1 = 160/10 = 16 Times

Company 2 = 79.20/8 = 9.9 Times

Company 3 = 78/6 = 13 Times

Company 4 = 297/45 = 6.6 Times